# Variational Search Distributions

### v2

@@ E.3 Approximation error for NN-based CPEs
$t\geq 1,\quad\hat{\sigma}_{t}^{2}(\mathbf{x})=k(\mathbf{x},\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+\rho\mathbf{I})^{-1}{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})\,,\quad\mathbf{x}\in\mathcal{X},$ (96)
which is equivalent to a GP posterior variance when observation noise is assumed to be zero-mean Gaussian with variance given by $\rho$. Given its similarities, we have that if enough sampling is asymptotically guaranteed, we can apply the same convergence results available for the GP-PI-based CPE, i.e., $\hat{\sigma}_{t}^{2}\in{\mathcal{{O}}}(t^{-1/2})$ almost surely. We then invoke E.5 to derive the following result.
Lemma E.1.
Let E.1 and E.5 hold. Then the following almost surely holds for the proxy variance:
@@
The proof follows by verifying that the sum of sampling probabilities at any point $\mathbf{x}\in\mathcal{X}$ diverges as $t\to\infty$ by E.5, and then by D.5 the result follows. ∎
Lemma E.2.
Let assumptions E.1 to E.4 hold. Then, given any $\delta\in(0,1]$, the following holds with probability at least $1-\delta$ for the approximation error between $\hat{\pi}_{t}$ and $\pi^{*}$:
@@
where $\beta_{t}(\delta):=\lVert\pi^{*}\rVert_{k}+\sigma_{\zeta}\sqrt{2\rho^{-1}\log(\det(\mathbf{I}+\rho^{-1}{\boldsymbol{\mathbf{{K}}}}_{t})^{1/2}/\delta)}$.
@@
The result above is a direct application of Theorem 3.5 in Maillard (2016) which provides an upper confidence bound on the kernelized least-squares regressor approximation error (another version of the same result is also available in Durand et al. (2018, Thr. 1)). ∎
@@
Proposition E.2.
Let assumptions E.1 to E.5 hold. Then, given $\delta\in(0,1]$, the following holds with probability at least $1-\delta$ for VSD equipped with a wide enough NN-based CPE model $\hat{\pi}_{t}$:
$\mathbb{D}\!\left[{p({\mathbf{x}}|{y>\tau_{t},\mathcal{D}_{t}})}\|{p({\mathbf{x}}|{y>\tau_{t},{f\!\centerdot}})}\right]\in{\mathcal{{O}}}\left(\sqrt{\frac{\xi_{t}}{t}}\right)\,.$ (100)
@@
The result follows by applying the same steps as in the proof of Sec. 2.4. We note that $\ell_{t}^{*}(\mathbf{x})=\pi^{*}(\mathbf{x})>0$, due to observation noise, so that $\ell_{t}^{*}(\mathbf{x})^{-1}\in{\mathcal{{O}}}_{\mathbb{P}}(1)$. Similarly, E.2 implies that $|\hat{\pi}_{t}(\mathbf{x})-\pi^{*}(\mathbf{x})|\leq\beta_{t}(\delta)\sigma_{t}(\mathbf{x})$ with probability at least $1-\delta$ simultaneously over all $\mathbf{x}\in\mathcal{X}$, so that ratio-dependent terms in Sec. 2.4 should remain bounded in probability. The upper bound in the result then follows by noticing that in our case $|{\Delta\ell}_{t}(\mathbf{x})|\leq\beta_{t}(\delta)\hat{\sigma}_{t}(\mathbf{x})$ with high probability, where $\hat{\sigma}_{t}\in{\mathcal{{O}}}(t^{-1/2})$ by E.1, and $\beta_{t}(\delta)\in{\mathcal{{O}}}(\sqrt{\xi_{t}})$ by E.2 and the definition of $\xi_{t}$ in Equation 99. ∎
The result above tells us that VSD is able to recover a similar asymptotic convergence guarantee to the one we derived for the GP-PI case, depending on the choice of NN architecture and more specifically on the spectrum of its associated NTK. In the case of a fully connected multi-layer ReLU network, for example, Chen & Xu (2021) showed an equivalence between the RKHS of the ReLU NTK and that of the Laplace kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-C\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert)$. As the latter is equivalent to a Matérn kernel with smoothness parameter set to $0.5$ (Rasmussen & Williams, 2006), the corresponding information gain bound is $\xi_{t}\in\widetilde{\mathcal{{O}}}(t^{\frac{d}{1+d}})$, where $d$ here denotes the dimensionality of the domain $\mathcal{X}$ (Vakili & Olkhovskaya, 2023). In the case of discrete sequences of length $M$, the dimensionality of $\mathcal{X}$ is determined by $M$. Therefore, in this case, we have proven Sec. 2.4.444Here $\widetilde{{\mathcal{{O}}}}_{\mathbb{P}}$ suppresses logarithmic factors, as in $\widetilde{{\mathcal{{O}}}}$.
@@
Appendix F VSD as a Black-Box Optimization Lower Bound

### v4

@@ F.3 Approximation error for NN-based CPEs
$t\geq 1,\quad\hat{\sigma}_{t}^{2}(\mathbf{x})=k(\mathbf{x},\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})\,,\quad\mathbf{x}\in\mathcal{X},$ (101)
where ${\boldsymbol{\mathbf{{\Sigma}}}}_{t}$ is the implicit regularization matrix in Equation 91 due to early stopping. The proxy variance is then equivalent to a GP posterior variance under the assumption of heteroscedastic (i.e., input dependent) Gaussian noise with covariance matrix given by ${\boldsymbol{\mathbf{{\Sigma}}}}_{t}$. Given its similarities, we have that if enough sampling is asymptotically guaranteed, we can apply the same convergence results available for the GP-PI-based CPE, i.e., $\hat{\sigma}_{t}^{2}\in{\mathcal{{O}}}(t^{-1})$ almost surely.
Lemma F.2.
Let assumptions F.2, F.5 and F.6 hold. Then the following almost surely holds for the proxy variance:
@@
We first observe that $\hat{\sigma}_{t}^{2}$ (101) is upper bounded by the posterior predictive variance of a GP model assuming i.i.d. Gaussian noise (cf. Sec. E.1) with variance $\rho^{*}$ satisfying:
$\rho^{*}\geq\lambda_{\max}({\boldsymbol{\mathbf{{\Sigma}}}}_{t})\,,\quad\forall t\in\mathbb{N}\,,$ (103)
which is such that:
$\begin{split}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}&={\boldsymbol{\mathbf{{K}}}}_{t}(e^{\nu_{t}s{\boldsymbol{\mathbf{{K}}}}_{t}}-\mathbf{I})^{-1}\\ &\preceq{\boldsymbol{\mathbf{{K}}}}_{t}(\nu_{t}s{\boldsymbol{\mathbf{{K}}}}_{t})^{-1}\\ &\preceq\frac{1}{s\nu_{t}}\mathbf{I}\end{split}$ (104)
since $e^{{\boldsymbol{\mathbf{{A}}}}}\succeq\mathbf{I}+{\boldsymbol{\mathbf{{A}}}}$, for any Hermitian matrix ${\boldsymbol{\mathbf{{A}}}}$, where $\succeq$ denotes the Loewner partial ordering in the space of positive-semidefinite matrices, i.e., ${\boldsymbol{\mathbf{{A}}}}\succeq{\boldsymbol{\mathbf{{B}}}}$ if and only if ${\boldsymbol{\mathbf{{A}}}}-{\boldsymbol{\mathbf{{B}}}}$ is positive semidefinite. Noting that the sum of sampling probabilities at any point $\mathbf{x}\in\mathcal{X}$ diverges as $t\to\infty$ by F.5, the result then follows by applying E.5 to the GP predictive variance upper bound with noise variance set to $\rho^{*}:=(s\nu_{*})^{-1}$ (F.6). ∎
Lemma F.3.
Let assumptions F.1 to F.6 hold. Then, given any $\delta\in(0,1]$, the following holds with probability at least $1-\delta$ for the approximation error between $\hat{\pi}_{t}$ and $\pi^{*}$:
@@
where $\beta_{t}(\delta):=\lVert\pi^{*}\rVert_{k}+\sigma_{\zeta}\sqrt{2\rho^{-1}\log(\det(\mathbf{I}+\rho^{-1}{\boldsymbol{\mathbf{{K}}}}_{t})^{1/2}/\delta)}$, and $\rho:=\frac{e^{-s}}{s\nu_{*}}$.
@@
The result above is a direct application of Theorem 3.5 in Maillard (2016) which provides an upper confidence bound on the kernelized least-squares regressor approximation error (another version of the same result is also available in Durand et al. (2018, Thr. 1)).
Let $\zeta_{i}:=z_{i}-\pi^{*}(\mathbf{x}_{i})$ denote the label noise in observation $i$, for $i\in\{1,\dots,t\}$. Expanding the definition of $\hat{\pi}_{t}$ (91) with $\pi_{0}=0$, given any $t\in\mathbb{N}$ and $\mathbf{x}\in\mathcal{X}$, we can decompose the approximation error as:
$\begin{split}|\pi^{*}(\mathbf{x})-\hat{\pi}_{t}(\mathbf{x})|&=|\pi^{*}(\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{z}}}_{t}|\\ &=|{\pi}^{*}(\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}({\boldsymbol{\mathbf{{\pi}}}}_{t}^{*}+{\boldsymbol{\mathbf{\zeta}}}_{t})|\\ &\leq|{\pi}^{*}(\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{{\pi}}}}_{t}^{*}|+|{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}|\end{split}$ (106)
where we applied the triangle inequality to obtain the last line. Analyzing the two terms on the right-hand side, by the reproducing property, we now have for the first term:
$\begin{split}|{\pi}^{*}(\mathbf{x})-{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{{\pi}}}}_{t}^{*}|&=|\langle\pi^{*},(\mathbf{I}-{\boldsymbol{\mathbf{{\Phi}}}}_{t}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})\varphi(\mathbf{x})\rangle_{k}|\\ &=|\langle\pi^{*},(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}\varphi(\mathbf{x})\rangle_{k}|\\ &\leq\lVert{\pi}^{*}\rVert_{k}\lVert(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}\varphi(\mathbf{x})\rVert_{k}\\ &=\lVert{\pi}^{*}\rVert_{k}\sqrt{\varphi(\mathbf{x})^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-2}\varphi(\mathbf{x})}\\ &\leq\lVert{\pi}^{*}\rVert_{k}\sqrt{\varphi(\mathbf{x})^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}\varphi(\mathbf{x})}\\ &=\lVert{\pi}^{*}\rVert_{k}\hat{\sigma}_{t}(\mathbf{x}),\end{split}$ (107)
where the second equality follows by an application of Woodbury's identity, the first inequality is due to Cauchy-Schwarz, the second inequality is due to the fact that ${\boldsymbol{\mathbf{{A}}}}^{-2}\preceq{\boldsymbol{\mathbf{{A}}}}^{-1}$ whenever ${\boldsymbol{\mathbf{{A}}}}\succeq\mathbf{I}$, and the last line follows from the definition of $\hat{\sigma}_{t}^{2}$. For the remaining in term (106), we have that:
$\begin{split}|{\boldsymbol{\mathbf{k}}}_{t}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}|&=|\langle\varphi(\mathbf{x}),{\boldsymbol{\mathbf{{\Phi}}}}_{t}({\boldsymbol{\mathbf{{K}}}}_{t}+{\boldsymbol{\mathbf{{\Sigma}}}}_{t})^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rangle_{k}|\\ &=|\langle\varphi(\mathbf{x}),(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rangle_{k}|\\ &=|\langle(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}\varphi(\mathbf{x}),(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rangle_{k}|\\ &\leq\sqrt{\varphi(\mathbf{x})^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}\varphi(\mathbf{x})}\lVert(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rVert_{k}\\ &=\hat{\sigma}_{t}(\mathbf{x})\sqrt{{\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}},\end{split}$ (108)
where we applied the identity ${\boldsymbol{\mathbf{{B}}}}^{\top}({\boldsymbol{\mathbf{{B}}}}{\boldsymbol{\mathbf{{B}}}}^{\top}+{\boldsymbol{\mathbf{{A}}}})^{-1}=(\mathbf{I}+{\boldsymbol{\mathbf{{B}}}}^{\top}{\boldsymbol{\mathbf{{A}}}}^{-1}{\boldsymbol{\mathbf{{B}}}})^{-1}{\boldsymbol{\mathbf{{B}}}}^{\top}{\boldsymbol{\mathbf{{A}}}}^{-1}$, which holds for an invertible matrix ${\boldsymbol{\mathbf{{A}}}}$ (Searle, 1982), to obtain the second equality, and the upper bound follows by the Cauchy-Schwarz inequality. For the norm of the noise-dependent term, we will apply a concentration bound by Abbasi-Yadkori (2012), which first requires a few transformations towards a non-time-varying regularization factor. Applying the SVD ${\boldsymbol{\mathbf{{\Phi}}}}_{t}={\boldsymbol{\mathbf{{U}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}$ (Mollenhauer et al., 2020), as in the proof of F.1, we have that ${\boldsymbol{\mathbf{{\Sigma}}}}_{t}={\boldsymbol{\mathbf{{V}}}}_{t}{\boldsymbol{\mathbf{{D}}}}_{t}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}$, where ${\boldsymbol{\mathbf{{D}}}}_{t}:={\boldsymbol{\mathbf{{\Lambda}}}}_{t}(e^{\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}}-\mathbf{I})^{-1}$ and ${\boldsymbol{\mathbf{{\Lambda}}}}_{t}={\boldsymbol{\mathbf{{S}}}}_{t}^{2}$, which leads us to:
$\begin{split}\lVert(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rVert_{k}^{2}&={\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &={\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}{\boldsymbol{\mathbf{{D}}}}_{t}^{-1}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{U}}}}_{t}^{\top}(\mathbf{I}+{\boldsymbol{\mathbf{{U}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{D}}}}_{t}^{-1}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{U}}}}_{t}^{\top})^{-1}{\boldsymbol{\mathbf{{U}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{D}}}}_{t}^{-1}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &={\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}^{2}{\boldsymbol{\mathbf{{D}}}}_{t}^{-2}(\mathbf{I}+{\boldsymbol{\mathbf{{D}}}}_{t}^{-1}{\boldsymbol{\mathbf{{S}}}}_{t}^{2})^{-1}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &={\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}({\boldsymbol{\mathbf{{S}}}}_{t}^{-2}{\boldsymbol{\mathbf{{D}}}}_{t}^{2}+{\boldsymbol{\mathbf{{D}}}}_{t})^{-1}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\,,\end{split}$ (109)
where we applied the identity $(\mathbf{I}+{\boldsymbol{\mathbf{{A}}}}{\boldsymbol{\mathbf{{B}}}})^{-1}{\boldsymbol{\mathbf{{A}}}}={\boldsymbol{\mathbf{{A}}}}(\mathbf{I}+{\boldsymbol{\mathbf{{B}}}}{\boldsymbol{\mathbf{{A}}}})^{-1}$ (Searle, 1982). For the eigenvalues of ${\boldsymbol{\mathbf{{\Sigma}}}}_{t}$, we have the following lower bound:
$\begin{split}{\boldsymbol{\mathbf{{D}}}}_{t}&={\boldsymbol{\mathbf{{\Lambda}}}}_{t}(e^{\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}}-\mathbf{I})^{-1}\\ &={\boldsymbol{\mathbf{{\Lambda}}}}_{t}e^{-\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}}(\mathbf{I}-e^{-\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}})^{-1}\\ &\succeq{\boldsymbol{\mathbf{{\Lambda}}}}_{t}e^{-\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}}(\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t})^{-1}\\ &=\frac{1}{\nu_{t}s}e^{-\nu_{t}s{\boldsymbol{\mathbf{{\Lambda}}}}_{t}}\\ &\succeq\frac{1}{\nu_{*}s}e^{-s}\mathbf{I}\,,\end{split}$ (110)
where the first inequality is due to $e^{-{\boldsymbol{\mathbf{{A}}}}}\succeq\mathbf{I}-{\boldsymbol{\mathbf{{A}}}}$, and the last inequality holds by F.6. Hence, setting $\rho:=\frac{e^{-s}}{s\nu_{*}}$, we have that:
$\begin{split}\lVert(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rVert_{k}^{2}&\leq{\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}(\rho^{2}{\boldsymbol{\mathbf{{S}}}}_{t}^{-2}+\rho\mathbf{I})^{-1}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &=\rho{\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}(\rho\mathbf{I}+{\boldsymbol{\mathbf{{S}}}}_{t}^{2})^{-1}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &=\rho^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{V}}}}_{t}{\boldsymbol{\mathbf{{S}}}}_{t}(\rho\mathbf{I}+{\boldsymbol{\mathbf{{S}}}}_{t}^{2})^{-1}{\boldsymbol{\mathbf{{S}}}}_{t}{\boldsymbol{\mathbf{{V}}}}_{t}^{\top}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &=\rho^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}^{\top}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top}(\rho\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{\zeta}}}_{t}\\ &=\lVert(\rho\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{\zeta}}}_{t}\rVert_{k}^{2}\end{split}$ (111)
By Abbasi-Yadkori (2012, Cor. 3.6), given any $\delta\in(0,1]$, we then have that the following holds with probability at least $1-\delta$:
$\forall t\geq 1,\quad\lVert(\mathbf{I}+{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{t}^{\top})^{-1/2}{\boldsymbol{\mathbf{{\Phi}}}}_{t}{\boldsymbol{\mathbf{{\Sigma}}}}_{t}^{-1}{\boldsymbol{\mathbf{\zeta}}}_{t}\rVert_{k}^{2}\leq 2\sigma_{\zeta}^{2}\log\left(\frac{\det(\mathbf{I}+\rho^{-1}{\boldsymbol{\mathbf{{K}}}}_{t})^{1/2}}{\delta}\right)\,.$ (112)
Finally, combining the bounds above into Equation 106 leads to the result in F.3. ∎
@@
Proposition F.2.
Let assumptions F.1 to F.6 hold. Then, given $\delta\in(0,1]$, the following holds with probability at least $1-\delta$ for VSD equipped with a wide enough NN-based CPE model $\hat{\pi}_{t}$:
$\mathbb{D}\!\left[{p({\mathbf{x}}|{y>\tau_{t},\mathcal{D}_{t}})}\|{p({\mathbf{x}}|{y>\tau_{t},{f\!\centerdot}})}\right]\in{\mathcal{{O}}}_{\mathbb{P}}\left(\sqrt{\frac{\xi_{t}}{t}}\right)\,.$ (114)
@@
The result follows by applying the same steps as in the proof of Sec. 2.4. We note that $\ell_{t}^{*}(\mathbf{x})=\pi^{*}(\mathbf{x})>0$, due to observation noise, so that $\ell_{t}^{*}(\mathbf{x})^{-1}\in{\mathcal{{O}}}_{\mathbb{P}}(1)$. Similarly, F.3 implies that $|\hat{\pi}_{t}(\mathbf{x})-\pi^{*}(\mathbf{x})|\leq\beta_{t}(\delta)\sigma_{t}(\mathbf{x})$ with probability at least $1-\delta$ simultaneously over all $\mathbf{x}\in\mathcal{X}$, so that ratio-dependent terms in Sec. 2.4 should remain bounded in probability. The upper bound in the result then follows by noticing that in our case $|{\Delta\ell}_{t}(\mathbf{x})|\leq\beta_{t}(\delta)\hat{\sigma}_{t}(\mathbf{x})$ with high probability, where $\hat{\sigma}_{t}\in{\mathcal{{O}}}(t^{-1/2})$ by F.2, and $\beta_{t}(\delta)\in{\mathcal{{O}}}(\sqrt{\xi_{t}})$ by F.3 and the definition of $\xi_{t}$ in Equation 113. ∎
The result above tells us that VSD equipped with an NN-based CPE can recover a similar asymptotic convergence guarantee to the one we derived for the GP-PI case, depending on the choice of NN architecture and more specifically on the spectrum of its associated NTK. In the case of a fully connected multi-layer ReLU network, for example, Chen & Xu (2021) showed an equivalence between the RKHS of the ReLU NTK and that of the Laplace kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-C\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert)$. As the latter is equivalent to a Matérn kernel with smoothness parameter set to $0.5$ (Rasmussen & Williams, 2006), the corresponding information gain bound is $\xi_{t}\in\widetilde{\mathcal{{O}}}(t^{\frac{d}{1+d}})$, where $d$ here denotes the dimensionality of the domain $\mathcal{X}$ (Vakili & Olkhovskaya, 2023). In the case of discrete sequences of length $M$, the dimensionality of $\mathcal{X}$ is determined by $M$. Hence, we have proven Sec. 2.4.444Here $\widetilde{{\mathcal{{O}}}}_{\mathbb{P}}$ suppresses logarithmic factors, as in $\widetilde{{\mathcal{{O}}}}$, and holds in probability.
@@
Appendix G VSD as a Black-Box Optimization Lower Bound